# Disentangling Memory and Reasoning Ability in Large Language Models

### v2

@@ 3.3.2 The impact of the number of special tokens
Another important issue is knowledge distillation. We need to ensure that the model’s improvement is not due to knowledge distillation from the GPT-4 framework. By using the same inference steps, we compared the results of standard training with 0 reason and memory tokens and found that adding these tokens significantly improves performance. This indirectly confirms that the model’s enhancement comes from algorithmic improvements rather than knowledge distillation.
3.4 Case Study
3.4.1 Sample Analysis
To validate the decoupling effect on memory and reasoning, we configure and employ GPT-4o-mini as an evaluator, assessing whether steps labeled as "memory" entail factual knowledge and those labeled as "reasoning" represent reasoning processes on StrategyQA, CommonSenseQA, and TruthfulQA. We use a structured, directive, one-shot Chain of Thought (CoT) prompting method as the baseline to compare with our method. The prompt setup is detailed in Appendix F.
In this CoT approach, directive prompting with explicitly instructs the model to distinguish memory information and reasoning steps. The one-shot example provides a structured format, demonstrating how memory (e.g., ) and reasoning parts (e.g., ) can be organized separately in answer generation (e.g., ). This structure guides the model to produce answers annotated as either memory or reasoning , enhancing interpretability by separating factual knowledge from reasoning processes.
From Figure 3 and Table 7, on the StrategyQA dataset, our method achieves an accuracy of 78.0% on LLaMA-3.1-8B, outperforming the One-shot CoT baseline by 20%, our approach achieves higher accuracy in decoupling memory (94% vs. 93%) and reasoning (71% vs. 67%), demonstrating effective decoupling between these two components in multi-steps inference. On the CommonsenseQA dataset, our method achieves an accuracy of 82.3% on LLaMA-3.1-8B, exceeding the One-shot CoT baseline by 26.3%. The results highlight that our approach consistently outperforms the baseline in decoupling memory (91% vs. 83%) and reasoning (78% vs. 74%), demonstrating robust performance in commonsense inference tasks. On the TruthfulQA dataset, our method achieves an accuracy of 86.6% on LLaMA-3.1-8B, surpassing the One-shot CoT baseline by 32.6%. The results further illustrate that our approach achieves superior accuracy in decoupling memory (89% vs. 88%) and reasoning (74% vs. 62%), highlighting its effectiveness in factual reasoning. Additionally, Table 7 shows that both LLaMA-3.1-8B and LLaMA-2-7B maintain consistent distributions of memory and reasoning across all datasets. This reflects the stability and generalizability of our decoupling mechanism, ensuring its applicability to diverse reasoning tasks.
@@ 3.4.1 Sample Analysis
| Method | Accuracy (% ) |
| StrategyQA | |
| One-shot CoT | 58.0 |
| Ours | 78.0 |
| CommonsenseQA | |
| One-shot CoT | 56.0 |
| Ours | 82.3 |
| TruthfulQA | |
| One-shot CoT | 54.0 |
| Ours | 86.6 |
@@
3.4.2 Attention Analysis
In the case study, we have found that the and do an important job in our LLM’s reasoning. Although using raw attention weights to interpret token importance can be somewhat controversial, attention patterns still provide valuable insights into how Transformers operate (Abnar & Zuidema, 2020). This heatmap, as Figure 4 shows that the model focuses intensely on specialized tokens throughout the CoT steps. These tokens received higher attention weights than regular tokens, suggesting they may play a more significant role in leading knowledge and reasoning content generation. This observation aligns with the main findings presented in the previous case study.
We input two sentences(which can be found in the appendix) into our fine-tuned LLaMA-3.1-8B model, resulting in a large attention heatmap. We then segmented the entire attention map according to the steps in the CoT, producing the two smaller maps below, as the large map was too cumbersome to analyze. Other attention maps can be found in the appendix. We observed that and tokens had relatively high attention values, indicating that the model places greater emphasis on this content, which indirectly demonstrates the effectiveness of our algorithm.
@@ 3.4.3 Error Analysis
For Example, the question 5 above emphasizes the role of the revolving door as a security measure, so the correct answer should be somewhat unexpected. Options B, C, and D represent typical uses of revolving doors for managing two-way traffic flow. Only in banks does a revolving door serve as a security measure. The correct answer is likely A. bank, as banks use revolving doors not only for easy access but also as a security measure to control entry and exit. The model’s knowledge is accurate, but this nuance was missed during reasoning.
@@ 4 Related Work
During pre-training, Large Language Models capture a large amount of knowledge in models’ parameters, known as parametric memory (Jin et al., 2024a). Previous research extensively explores the mechanism of inference with parametric memory, they observe that models can well adopt memory for simple tasks but struggle for complex inference, e.g., multi-hop inference (Li et al., 2024b; Yang et al., 2024b; Wang et al., 2024a). Others reveal the challenges in the leverage of parametric knowledge, particularly when dealing with long-tail facts (facts associated with less common entities) or when the knowledge is rare (Wang et al., 2023a; Allen-Zhu & Li, 2023; Cheng et al., 2024). These studies primarily focus on the analysis of model behavior. While valuable, they do not address how to better elicit the parametric knowledge for inference. In this work, we explore how to boost LLMs’ leverage of their parametric knowledge for complex inference.
@@
Recent research on enhancing large language models’ inference capabilities can be broadly categorized into prompt-based and tuning-based approaches. Prompt-based methods strategically guide reasoning processes. Chain-of-Thought (CoT) prompting (Wei et al., 2022b) and its derivatives (Zhao et al., 2024; Zhou et al., 2023; Chen et al., 2024a; Hu et al., 2023; Jin et al., 2024b) decompose complex tasks into sequential steps, improving transparency and decision-making. Others like Tree-of-Thoughts (ToT) (Yao et al., 2023) and Graph-of-Thoughts (GoT) (Besta et al., 2024) further utilize hierarchical and network-based inference to cover larger searching space. These strategies design a framework where LLMs can elicit the parametric memory relevant to the task. However, in these methods, the models might not know when to reason or use their memory. Tuning-based methods introduce trainable tokens for structured CoT steps, facilitating reasoning and utilization of memory Wang et al. (2023b); Goyal et al. (2023); Colon-Hernandez et al. (2024) Despite the effectiveness, these methods intertwine reasoning and memory usage, which may limit the full potential of the models. In contrast, our approach aims to decouple memory and reasoning within the CoT process by introducing various special tokens, enabling the model to leverage its memory more effectively.
@@ 5 Conclusion
In this work, we introduced a novel framework for training an autoregressive language model to distinguish between reasoning and memory processes. By utilizing two specific tokens— for retrieved factual knowledge and for reasoning steps—the model can effectively disentangle knowledge retrieval from logical inference, enhancing interpretability and performance on complex reasoning tasks. The structured input encoding enables the model to maintain a clear boundary between memory and reasoning, improving its generalization capabilities for queries requiring a blend of factual knowledge and multi-step reasoning. Through this approach, the model not only achieves more accurate answers but also provides interpretable outputs that clarify the reasoning process step-by-step, a feature critical for tasks in fields requiring transparency and accountability in AI-generated reasoning. Despite the promising results, our approach has certain limitations: As the number of reasoning steps grows, the input sequence length can become substantially large, especially for complex questions requiring multiple reasoning steps. This may lead to increased computational costs and memory limitations, especially when training large models. The model’s interpretability and performance improvements are closely tied to the effectiveness of the and annotations. Incorrect tagging or inconsistent use of these tokens during training could result in ambiguous or incorrect responses during inference.
@@ 6 Limitation
The proposed decomposition framework offers a promising approach to disentangling memory recall and reasoning in large language models. However, like any novel framework, its limitations should be acknowledged for future improvement. One limitation is the framework’s reliance on the quality and breadth of training data for memory recall. While this is a common issue for most machine learning models, missing or underrepresented knowledge in the training data could affect the accuracy and completeness of retrieval, particularly in niche domains. However, this limitation can be mitigated by incorporating dynamic updating mechanisms or integrating external knowledge bases to fill in gaps.
Introducing special tokens, such as and , while useful for distinguishing between memory retrieval and reasoning tasks, adds some complexity to the tokenization process. This complexity may require additional task-specific tuning and adaptation to different LLM architectures or language domains. Nevertheless, the token-based approach provides a clear scaffold for interpretability, which could offset the initial complexity by enhancing model transparency. The framework also faces challenges with tasks that require highly complex reasoning, such as multi-hop or deeply nested logic, which may not be easily separable into discrete memory retrieval and reasoning steps. While this suggests the need for further refinement in handling complex chains of thought, the framework still performs well for a wide range of tasks, demonstrating its robustness in standard reasoning scenarios.
Appendix A Appendix
SUMMARY OF THE APPENDIX
This appendix contains additional details for the “Disentangling Memory and Reasoning Ability in Large Language Models”. The appendix is organized as follows:
Appendix B Data Generation
B.1 Implement Details
@@
We developed an LLM-based data generation framework based on GPT-4o to generate high-quality training data for memory and reasoning steps. This framework includes two LLMs: an inference LLM and a knowledge LLM. The inference LLM is responsible for generating Chain-of-Thought (CoT) processes, decoupling memory from reasoning, and assigning labels to each sub-step by marking those requiring factual knowledge as [memory] and those requiring reasoning as [reason]. The prompt of the inference agent is shown as prompt 2 below. The Knowledge LLM retrieves the necessary knowledge for [memory] steps based on questions provided by the inference LLM. The prompt of the Knowledge LMM is shown as prompt 1 below. We use these two LLMs to ensure the independence of memory and reasoning within the CoT, providing high-quality data for subsequent training. Figure 6 is the prompt configuration for a knowledge LLM. The questions corresponds to the Knowledge base content in the inference LLM and is used to supply accurate factual information for steps labeled as <memory>.
@@
The prompt configuration for inference LLM in a zero-shot setting is shown in Figure 7. The Step name refers to the specific name of each step in the Chain of Thought (CoT) process. The Requirement column labels whether each step pertains to <memory> or <reason>. The Knowledge based is used to provide questions related to factual knowledge in <memory> steps, while the Content focuses on designing to outline the reasoning process for <reason> steps. This structure facilitates a clear distinction between memory retrieval and reasoning tasks, enhancing the model’s capability to execute complex sequences in a zero-shot environment.
@@ B.2 Example
In this study, we leverage both the inference LLM and knowledge LLM based on GPT-4o to generate a dataset. Specifically, we use the StrategyQA dataset as the source for our generation. The StrategyQA dataset is a question-answering dataset designed to evaluate models’ multi-hop inference abilities. It includes questions that require strategic thinking and often demand more than one piece of information to answer correctly.
Appendix C Preliminary
C.1 Experiment
C.2 Example
A value greater than or equal to 0.2 indicates that the two contents are very unrelated, like example 9.
Appendix D Experiment Details
D.1 Dataset
StrategyQA (Geva et al., 2021)
StrategyQA is a challenging question-answering benchmark that focuses on implicit, multi-step reasoning. Unlike conventional multi-hop datasets where questions explicitly outline the steps needed to reach an answer, StrategyQA requires models to infer these reasoning steps. Each question in StrategyQA is crafted to be implicit and short, with Boolean ("Yes" or "No") answers, requiring logical deductions based on general knowledge. For example, answering a question like “Did Aristotle use a laptop?” involves reasoning about the historical timeline of both Aristotle’s life and the invention of laptops. The StrategyQA dataset includes a total of 2,780 verified questions. The training set comprises 1,600 questions which are used for fine-tuning, and the validation (test) set contains 690 questions, which are used for the validation of each baseline and our method in our experiment.
CommonsenseQA (Talmor et al., 2019)
CommonsenseQA is a multiple-choice dataset with 12,247 questions aimed at testing AI on commonsense reasoning using the ConceptNet knowledge graph. Each question has one correct answer and four distractors, requiring models to understand relations like causality and spatial proximity. While humans achieve 88.9% accuracy, advanced models like BERT-Large reach only 55.9%, underscoring the challenge of commonsense reasoning in AI. The training set comprises 9,740 questions, the validation set contains 1,220 questions, which are used for fine-tuning in our experiment, and the tests set includes 1,140 questions, which are used for the validation of each baseline and our method in our experiment.
TruthfulQA (Lin et al., 2022)
@@ D.1 Dataset
D.2 Evaluation Metric
To mitigate the inherent output instability of LLMs in both CoT and Zero-shot settings, we found that conventional answer-matching techniques, such as regular expression-based methods, may not reliably capture the precise answers required. Consequently, we adopted GPT-4o-mini as an evaluation tool to compute the LLM performance across multiple datasets (Cao et al., 2024). This approach enables a more nuanced assessment of LLM outputs, given the limitations of regular matching techniques under these settings. The detailed prompt used for evaluation is shown in Prompt 11 below.
Appendix E Training Details
E.1 Training Configuration
@@ E.3 Example
Appendix F Case Study
F.1 Sample Analysis
For sample analysis, to highlight the decoupling effectiveness, reasoning capability, and interpretability of our approach, we set One-shot Chain-of-Thought (CoT) reasoning as the baseline for this evaluation (see details in Figure). We leverage GPT-4o-mini as an evaluator with the prompt configuration provided below to assess the decoupling effectiveness of LLaMA-3.1-8B in separating memory from reasoning processes.
@@ F.2 Error Analysis
F.3 More Attention Maps
@@
F.4 More Analysis Samples

### v3

@@ 3.3.2 Impact of Special Token Count
3.4 Further Analysis
In this section, we aim to analyze the decoupling effect 3.4.1, attention analysis 3.4.2 of our method and error analysis of our method for 3.4.3.
3.4.1 Decoupling Analysis
To validate the decoupling effect on memory and reasoning, we configure GPT-4o-mini as an evaluator (details in Appendix E.1), assessing whether steps labeled as "memory" entail factual knowledge and those labeled as "reasoning" represent reasoning processes on our three benchmarks. Then We use a structured, directive one-shot Chain of Thought (CoT) prompting method to prompt LLaMA-3.1-8B as the baseline that can also disentangle memory and reason step. This prompt setup is displayed in Appendix E.1 in Figure 20.
| Method | Accuracy (% ) |
| StrategyQA | |
| One-shot CoT | 58.0 |
| Ours | 78.0 |
| CommonsenseQA | |
| One-shot CoT | 56.0 |
| Ours | 82.3 |
| TruthfulQA | |
| One-shot CoT | 54.0 |
| Ours | 86.6 |
In this CoT approach, directive prompting with explicitly instructs the model to distinguish memory information and reasoning steps. The one-shot example provides a structured format, demonstrating how memory (e.g., ) and reasoning parts (e.g., ) can be organized separately in answer generation (e.g., ). This structure guides the model to produce answer and inference steps annotated as either memory or reasoning , enhancing interpretability by separating factual knowledge and reasoning processes.
From Figure 2 and Table 5, on the StrategyQA dataset, our method achieves an accuracy of 78.0% on LLaMA-3.1-8B, outperforming the One-shot CoT baseline by 20%, our approach achieves higher accuracy in decoupling memory (94% vs. 93%) and reasoning (71% vs. 67%), demonstrating effective decoupling between these two components in multi-steps inference. On the CommonsenseQA dataset, our method achieves an accuracy of 82.3% on LLaMA-3.1-8B, exceeding the One-shot CoT baseline by 26.3%. The results highlight that our approach consistently outperforms the baseline in decoupling memory (91% vs. 83%) and reasoning (78% vs. 74%), demonstrating robust performance in commonsense inference tasks. On the TruthfulQA dataset, our method achieves an accuracy of 86.6% on LLaMA-3.1-8B, surpassing the One-shot CoT baseline by 32.6%. The results further illustrate that our approach achieves superior accuracy in decoupling memory (89% vs. 88%) and reasoning (74% vs. 62%), highlighting its effectiveness in factual reasoning. Additionally, Table 6 shows that both LLaMA-3.1-8B and LLaMA-2-7B maintain consistent distributions of memory and reasoning across all datasets. This reflects the stability and generalizability of our decoupling mechanism, ensuring its applicability to diverse inference tasks.
@@ 3.4.1 Decoupling Analysis
3.4.2 Attention Analysis
In the case study 3.4.1, we have found that the and do an important job in our LLM’s Inference. Although using raw attention weights to interpret token importance can be somewhat controversial, attention patterns still provide valuable insights about how transformers operate Abnar and Zuidema (2020). This heatmap, as Figure 3 shows that the model focuses intensely on specialized tokens throughout the inference. These tokens received higher attention weights than regular tokens, suggesting they play a more significant role in leading knowledge and reasoning content generation. This observation aligns with the main findings presented in the previous case study 3.4.1.
We input two sentences (can be found in Appendix E.3 in Figure 18) into our fine-tuned LLaMA-3.1-8B model, getting a large attention heatmap. We then segmented two entire attention maps according to the steps by model inference, producing the two smaller maps above as Figure 3. Other samples can be found in the Appendix E.3. by the observation, it indicate that the model places greater emphasis on this content, which indirectly demonstrates the effectiveness of our algorithm.
@@ 3.4.3 Error Analysis
For Example, the question 4 above emphasizes the role of the revolving door as a security measure, so the correct answer should be somewhat unexpected. Options B, C, and D all represent typical uses of revolving doors for managing two-way traffic flow. Only in banks does a revolving door serve as a security measure. The correct answer is likely A. bank, as banks use revolving doors not only for easy access but also as a security measure to control entry and exit. The model’s knowledge is accurate, but it missed this nuance during reasoning steps.
@@ 4 Related Work
During pre-training, large language models capture a large amount of knowledge in the models’ parameters, known as parametric memory. Previous research extensively explores the mechanism of inference with parametric memory; they observe that models can well adopt memory for simple tasks but struggle with complex inference, e.g., multi-hop inference (Li et al., 2024b; Yang et al., 2024b; Wang et al., 2024a; Jin et al., 2025b; Han et al., 2024; Wang et al., 2025a). Others reveal the challenges in the leverage of parametric knowledge, particularly when dealing with long-tail facts (facts associated with less common entities) or when the knowledge is rare (Wang et al., 2023; Allen-Zhu and Li, ; Cheng et al., 2024; Jin et al., 2025a). These studies primarily focus on the analysis of model behavior. While valuable, they do not address how to better elicit the parametric knowledge for inference. In this work, we explore how to boost LLMs’ leverage of their parametric knowledge for complex inference.
@@
Recent research on enhancing LLMs’ inference capabilities can be broadly categorized into prompt-based and tuning-based approaches. Prompt-based methods strategically guide reasoning processes. Chain-of-Thought (CoT) prompting Wei et al. (2022b) and its derivatives Zhao et al. (2024); Zhou et al. (2023); Chen et al. (2024a); Hu et al. (2023); Jin et al. (2024) decompose complex tasks into sequential steps, improving transparency and decision-making. Others like Tree-of-Thoughts (ToT) Yao et al. (2023) and Graph-of-Thoughts (GoT) Besta et al. (2024) further utilize hierarchical and network-based inference to cover a larger search space. These strategies design a framework where LLMs can elicit the parametric memory relevant to the task. However, in these methods, the models might not know when to reason or use their memory.
Tuning-based methods introduce trainable tokens for structured CoT steps, facilitating reasoning and utilization of memory Wang et al. (2024d); Goyal et al. (2024); Colon-Hernandez et al. (2024); Wang et al. (2025b). Despite the effectiveness, these methods intertwine reasoning and memory usage, which may limit the full potential of the models. In contrast, our approach aims to decouple memory and reasoning within the CoT process by introducing various special tokens, enabling the model to leverage its memory more effectively.
@@ 5 Conclusion
In this work, we proposed a novel inference framework for training LLMs to distinguish between reasoning and memory processes using two special tokens: for factual knowledge retrieval and for logical reasoning. This structured input disentangles these processes, enhancing interpretability and improving performance on complex reasoning tasks. By maintaining a clear boundary between memory and reasoning during training, the model generalizes better to queries that combine factual knowledge with multi-step reasoning. This approach not only ensures more accurate answers but also produces interpretable, step-by-step reasoning outputs, crucial for transparency and accountability in complex reasoning.
@@ 6 Limitation
The proposed decomposition framework provides a promising method to disentangle memory recall and reasoning in large language models, enhancing interpretability and modularity. However, it has limitations that offer opportunities for improvement. One challenge is its reliance on the quality and breadth of training data for memory recall, which may lead to incomplete retrieval in underrepresented domains. This issue, common in machine learning, can be mitigated through dynamic updates or integration with external knowledge bases. The use of special tokens like and simplifies distinguishing between tasks but adds complexity to tokenization, requiring task-specific tuning for different architectures or languages. Nonetheless, this token-based design enhances transparency, offsetting the added complexity. The framework also struggles with tasks requiring deeply nested or multi-hop reasoning, as these steps may not neatly separate into recall and reasoning phases. Further refinement is needed to better handle complex reasoning chains, though the framework performs robustly in standard scenarios. Additionally, the retrieval-based approach introduces computational overhead, which may limit real-time applicability. However, the trade-off for interpretability and error traceability is valuable for use cases where transparency is critical, making this framework a significant step forward for addressing reasoning and memory in LLMs
Appendix A Data Generation
A.1 Implement Details
@@
We developed an LLM-based data generation framework based on GPT-4o to generate high-quality training data for decoupling memory and reasoning steps. This framework includes two LLMs: an inference LLM and a knowledge LLM. The inference LLM is responsible for generating Chain-of-Thought (CoT) inference processes, decoupling memory and reasoning, and then further assigning labels to each sub-step by marking those requiring factual knowledge as [memory] and those requiring reasoning as [reason]. The prompt of the knowledge agent is shown in Figure 5. The Knowledge LLM retrieves the necessary knowledge for [memory] steps based on questions provided by the inference LLM. We use these two LLMs to ensure the independence of memory and reasoning within the CoT, providing high-quality data for subsequent training. Figure 6 is the prompt configuration for inference LLM. The questions corresponds to the Knowledge base content in the inference LLM and is used to supply accurate factual information for steps labeled as <memory>.
@@
The Step name refers to the specific name of each step in the Chain of Thought (CoT) process. The Requirement labels whether each step pertains to <memory> or <reason>. The Knowledge based is used to provide questions related to factual knowledge in <memory> steps, while the Content focuses on designing to outline the reasoning process for <reason> steps. This structure facilitates a clear distinction between memory retrieval and reasoning tasks, enhancing the model’s capability to execute complex sequences in a zero-shot environment.
A.2 Example
In this study, we leverage both the inference LLM and knowledge LLM based on GPT-4o to generate a dataset. Specifically, we use the StrategyQA dataset as the source for our generation. The StrategyQA dataset is a question-answering dataset designed to evaluate models’ multi-hop reasoning abilities. It includes questions that require strategic thinking and often demand more than one piece of information to answer correctly.
Appendix B Preliminary
B.1 Experiment
@@ B.2 Example
A value greater than or equal to 0.2 indicates that the two contents are very unrelated like example 8.
Appendix C Experiment Details
C.1 Dataset
StrategyQA Geva et al. (2021)
StrategyQA is a challenging question-answering benchmark that focuses on implicit, multi-step reasoning. Unlike conventional multi-hop datasets where questions explicitly outline the steps needed to reach an answer, StrategyQA requires models to infer these reasoning steps. Each question in StrategyQA is crafted to be implicit and short, with Boolean ("Yes" or "No") answers, requiring logical deductions based on general knowledge. For example, answering a question like “Did Aristotle use a laptop?” involves reasoning about the historical timeline of both Aristotle’s life and the invention of laptops. The StrategyQA dataset includes a total of 2,780 verified questions. The training set comprises 1,600 questions which are used for fine-tuning, and the validation (test) set contains 690 questions, which are used for the validation of baselines and our method in our experiment.
CommonsenseQA Talmor et al. (2019)
CommonsenseQA is a multiple-choice dataset with 12,247 questions aimed at testing AI on commonsense reasoning using the ConceptNet knowledge graph. Each question has one correct answer and four distractors, requiring models to understand relations like causality and spatial proximity. While humans achieve 88.9% accuracy, advanced models like BERT-Large reach only 55.9%, underscoring the challenge of commonsense inference in AI. The training set comprises 9,740 questions, the validation set contains 1,220 questions, which are used for fine-tuning, and the test set includes 1,140 questions, which are used for the validation of baselines and our method in our experiment.
TruthfulQA Lin et al. (2022)
@@ C.1 Dataset
C.2 Evaluation Metric
To mitigate the inherent output instability of LLMs in both CoT and Zero-shot settings, we found that conventional answer-matching techniques, such as regular expression-based methods, may not reliably capture the precise answers required. Consequently, we adopted GPT-4o-mini as an evaluation tool to compute the LLM performance across multiple datasets Cao et al. (2024). This approach enables a more nuanced assessment of LLM outputs, given the limitations of regular matching techniques under these settings. The detailed prompt used for evaluation is shown in Prompt 10 below.
Appendix D Training Details
D.1 Training Configuration
@@ D.3 Example
Appendix E Case Study
E.1 Sample Analysis
For sample analysis, to highlight the decoupling effectiveness, reasoning capability, and interpretability of our approach, we set One-shot Chain-of-Thought (CoT) reasoning as the baseline for this evaluation, see details in Figure 20. We leverage GPT-4o-mini as an evaluator with prompt configuration provided as below to assess the decoupling effectiveness of LLaMA-3.1-8B in separating memory and reasoning processes.
@@ E.2 Error Analysis
E.3 More Attention Maps
@@
E.4 More Analysis Samples